Calculate diffusion over potential for hydrogen evolution reaction, 2H^+ + 2e- = H2, at a current density of .2 A/cm2 and temperature 25C if limiting current density is 0.5 A/cm2? beta =.5
Answer #1

Diffusion rate \alpha 1/\sqrt{D} D = Density

D1 = 0.2 and D2 = 0.5 r1 , r2 are the rate of diffusion

r1/r2 =\sqrt{D2 /D1}

r1/r2 = \sqrt{0.5 /0.2}

r1/r2 =\sqrt{2.5}

r1 /r2 = 1.58

r1 = 1.58 r2